OpenStudy (kainui):

I'm looking for a closed form for this, not sure if it exists:

OpenStudy (kainui):

\[f(x,a)=\sum_{n=0}^x \frac{(-a)^n}{n!}\]

OpenStudy (kainui):

Basically the way I was thinking of handling it was to just write it as \[f(x,a)=\sum_{n=0}^x \frac{(-a)^n}{n!}=e^{-a}-\sum_{n=x+1}^\infty \frac{(-a)^n}{n!}\] and then say for large x, the sum can be thrown away but I don't like that very much.

imqwerty (imqwerty):

closed form?

OpenStudy (kainui):

Yeah, so if I plug in a value of x, I don't want to have to add up all the terms form n=0 to n=x haha.

ganeshie8 (ganeshie8):

you're looking for a more closed form of \[\Gamma(x,a)=(x-1)!e^{-a}\sum\limits_{n=0}^{x-1}\dfrac{a^k}{n!}\] ?

ganeshie8 (ganeshie8):

\[\Gamma(a,x) = \int\limits_{x}^{\infty} t^{a-1}e^{-t}\,dt\]

ganeshie8 (ganeshie8):

not sure how that is any better than the partial sum that you have...

OpenStudy (ikram002p):

what do u mean by from n=0 to x ? what is x ?

OpenStudy (kainui):

I don't think that's what I'm looking for, I guess I should clarify that I am mostly interested in x being an integer greater than 0.

OpenStudy (ikram002p):

and a can be anything ?

OpenStudy (ikram002p):

are u messing for Fourier ?

OpenStudy (kainui):

\[f(x,a)=\sum_{n=0}^x \frac{(-a)^n}{n!}\] So for example: \[f(3,a)=\sum_{n=0}^3 \frac{(-a)^n}{n!} = 1 - a + \frac{a^2}{2}-\frac{a^3}{6}\] a can be anything, real or complex, but I think it's best to keep it to positive integer for now but it really doesn't matter.

OpenStudy (kainui):

Specifically I'm looking at solving this recurrence relation: \[f(x+1)+(x+1)f(x) =a^x\]

OpenStudy (kainui):

Oh well, I already found the solution and it's not too difficult to plug it in and see that it does indeed satisfy the equation. \[f(x)=(-1)^x x! \left(f(0) + \sum_{n=0}^x \frac{(-a)^n}{n!} \right)\]

ganeshie8 (ganeshie8):

\(f(x) = a^{x-1} - xf(x-1) \\~\\= a^{x-1} - x(a^{x-2}-(x-1)f(x-2))\\~\\ =a^{x-1}-xa^{x-2}+x(x-1)f(x-2)\\~\\ = \cdots \)

ganeshie8 (ganeshie8):

Okay I think I see how you got that weird sum :)

OpenStudy (kainui):

Yeah I discovered that for recurrence relations of this form with p(x), q(x) and f(0) given: $$f(x+1)+p(x)f(x)=q(x)$$ Has solution: $$f(x)=(-1)^x\prod_{n<x} p(n) \left( f(0)+\sum_{n=0}^{x-1}\frac{(-1)^{n+1}}{\prod_{k \le n} p(k)}q(n) \right)$$

OpenStudy (kainui):

Anyways I'm just sorta trying to test it out on random stuff to make sure it works, kinda weird and fun haha. The problem is I don't know many closed form functions that look like these, which is why I picked specifically \(p(x)=x+1\) and \(q(x)=a^x\) because it's what we started with here